# Adaptive Exploration for Data-Efficient General Value Function Evaluations

### v1

@@ 5 Theoretical Framework
The proof is presented in Appendix A. ∎
Theorem 5.1 outlined above provides an iterative solution using the variance under the previous behavior policy $\mu_{k}$ to guide the next policy $\mu_{k+1}$. These updates drive the behavior policy to explore areas with higher variability, leading to a decrease in the overall variance and, hence, MSE. Furthermore, this approach of reducing overall variance also allows accurate learning of multiple GVF values with fewer data samples, thereby reducing the number of environmental interactions needed to obtain good approximations. In the next theorem, we will establish that each update from $\mu_{k}$ to $\mu_{k+1}$ decrease the overall variance across GVFs, providing monotonic progress towards the overall objective in an oscillation-free manner.
@@ 6 Variance Function
The theorems given in the previous section suggest an approach to learn an optimal behavior policy is dependent on the variance function $M_{\pi_{i}}^{\mu_{k}}$. We will study this variance function in detail now.
@@
In an off-policy context, Jain et al. (2021) introduced the variance function $M$, which estimates the variance in return under a target policy $\pi$ using data from a different behavior policy $\mu$. We will directly use this function $M$ as our variance estimator and present it here for completeness. The function $M$ for a state-action pair under $\pi$, with an importance sampling correction factor $\rho_{t}=\frac{\pi(a_{t}|s_{t})}{\mu(a_{t}|s_{t})}$, is defined as:
$M_{\pi}^{\mu}(s,a)$ $=\text{Var}_{a\sim\mu}\left(G_{t,\pi}|s_{t}=s,a_{t}=a\right)=\mathbb{E}_{a\sim\mu}\left[\left(G_{t,\pi}-\mathbb{E}_{a\sim\mu}[G_{t,\pi}|s_{t}=s,a_{t}=a]\right)^{2}|s_{t}=s,a_{t}=a\right]$ (5)
$=\mathbb{E}_{a\sim\mu}\left[\delta_{t}^{2}+\gamma^{2}\rho_{t+1}^{2}M_{\pi}^{\mu}(s_{t+1},a_{t+1})|s_{t}=s,a_{t}=a\right]$
Here, $G_{t,\pi}$ is the return at time $t$, and $\delta_{t}=r_{t}+\gamma V_{\pi}(s_{t+1})-V_{\pi}(s_{t})$ (Sutton & Barto, 2018) is the TD error. The variance function $M_{\pi}^{\mu}(s,a)$ relates the variance under $\pi$ from the current state-action pair to the next, when trajectories are sampled from $\mu$. This allows us to effectively bootstrap and iteratively update using temporal difference (TD) style method. Similarly, variance given state:
$M_{\pi}^{\mu}(s)$ $=\text{Var}_{a\sim\mu}\left(G_{t,\pi}|s_{t}=s\right)=\mathbb{E}_{a\sim\mu}\left[\rho_{t}^{2}\left(\delta_{t}^{2}+\gamma^{2}\rho_{t+1}^{2}M_{\pi}^{\mu}(s_{t+1},a_{t+1})\right)|s_{t}=s\right]$ (6)
Both state and state-action variances are interconnected, with $M^{\mu}_{\pi}(s)=\sum_{a}\mu(a|s)\rho^{2}(s,a)M^{\mu}_{\pi}(s,a)$. Note that the value function $V_{\pi}$ is required to compute the TD error $\delta$. Following Jain et al. (2021), we substitute the value estimate $\hat{V}$ for the true function $V_{\pi}$ to compute $\delta_{t}$ in Eq.(5) and Eq.(6).
Next, we prove the existence of $M$, which was not actually covered in Jain et al. (2021). This existence proof inherently establishes an upper bound on the IS ratio $\rho$. This condition limits the divergence of the behavior policy from the target policy, aligning with methodologies like TRPO (Schulman et al., 2015) and Retrace (Munos et al., 2016), which control the stability of policy updates by regulating the divergence.
When does the Variance Function Exists?
@@
(Variance Function $M$ Existence:) Given a discount factor $0<\gamma\leq 1$, the existence of variance function $M$ depends on the invertibility of matrix $(I-\gamma^{2}\mathbf{\bar{P}_{\mu}})$ if the below condition is satisfied:
$\mathbb{E}_{a\sim\mu}\left[\rho^{2}(s,a)\right]<\frac{1}{\gamma^{2}},\forall{s\in\mathcal{S}}.$
@@
Proof in Appendix A. ∎
Lemma 6.1 results in a loose upper bound on the IS weight $\rho$. This constraint emerges naturally as a necessary condition for the existence of the variance operator in Equation 7.
@@ 7 Algorithm
We are now ready to discuss the GVFExplorer algorithm, detailed in Algorithm 1. Our approach uses two networks: one for the value function ($Q$) and another for the variance function ($M$), each with $\mathcal{N}$ heads (one head for each GVF). The agent begins with a randomly initialized behavior policy. At each step, the agent observes cumulants for each of the $\mathcal{N}$ GVFs, then it updates the $Q$-network using off-policy TD. In our experiments, we use Expected Sarsa (Sutton & Barto, 2018), which has lower variance than other off-policy learning methods. This happens because Expected Sarsa updates the target with the expected value over the target policy’s actions for the next state value. The TD-error $\delta$ generated by the $Q$-network is then used to update the variance network $M$ (see Equation 5). We use Expected Sarsa for the $M$-network update as well. The behavior policy is then updated based on the new variance estimates, using Equation 4. This iterative process continues through $K$ environment steps, ultimately producing a refined $Q$-network for GVF value predictions.
To further enhance the algorithm’s adaptability and efficiency, we incorporate established Deep RL techniques such as Experience Replay Buffer for data reuse and target networks for both $Q$ and $M$ to improve learning stability.
Input: Target policies $\pi_{i\in\{1,2,\dots n\}}$, randomly initialize behavior policy $\mu_{1}$, replay buffer $\mathcal{D}$, randomly initialized primary networks $Q_{\theta},M_{w}$, and target networks $Q_{\bar{\theta}},M_{\bar{w}}$, learning rate for value $\alpha_{Q}$ and variance $\alpha_{M}$, mini-batch size $b$, trajectory length $T$, target network update frequency $l=100$, value and variance network update frequency $p=4,m=8$ respectively, training steps $K$
@@
2for environment step $=1,\dots K$ do
3       Observe state $s_{t}$, action $a_{t}\sim\mu_{k}(\cdot|s_{t})$, next state $s_{t+1}$, cumulants $c_{t}=\textit{Vector(size}(n))$
4       Store transition $(s_{t},a_{t},s_{t+1},c_{t})$ in replay $\mathcal{D}$.
@@
6             //Update the $Q_{\theta}$ network
@@
8             Compute $Q_{tar}$ as Tensor(size($b,n,|\mathcal{A}|$)).
9             Update $Q_{\theta}$ using gradient descent on MSE loss $(Q_{tar}(s_{t},a_{t})-Q_{\theta}(s_{t},a_{t}))^{2}$.
10             if $step\%l==0$ then
11                   $\bar{\theta}=\theta$  //Update the target Q network weights $\bar{\theta}$
@@ 18             Sample mini-batch of size $b$ of transition $(s_{t},a_{t},s_{t+1},c_{t})\sim\mathcal{D}$
19             Compute target $M_{tar}$ as Tensor(size($b,n,|\mathcal{A}|$)), where,
$\delta_{t}=c_{t}+\gamma$ $\mathbb{E}_{a^{\prime}\sim\pi(s_{t+1})}\left[Q_{\bar{\theta}}(s_{t+1},a^{\prime})\right]-Q_{\theta}(s_{t},a_{t})$
$M_{tar}(s_{t},a_{t})$ $=\delta_{t}^{2}+\gamma^{2}\mathbb{E}_{a^{\prime}\sim\pi(s_{t+1})}\left[M_{\bar{w}}(s_{t+1},a^{\prime})\right]$
Update $M_{w}$ using gradient descent on MSE loss $(M_{tar}(s_{t},a_{t})-M_{w}(s_{t},a_{t}))^{2}$
@@
21                   $\bar{w}=w$  //Update target M network weights $\bar{w}$
22
23             end if
@@ 21                   $\bar{w}=w$  //Update target M network weights $\bar{w}$
Returns GVFs Values $\{Q_{\theta}\}_{i\in\{1,2\dots n\}}$.
@@ 8 Experiments
We investigate the empirical utility of our proposed algorithm in both discrete and continuous state environments. Our experiments are designed to answer the following questions: (a) can GVFExplorer evaluate multiple GVFs efficiently? (b) how does GVFExplorer compare with the different baselines (explained below) in terms of convergence speed and estimation quality? (c) can GVFExplorer handle a large number of GVFs? (d) can GVFExplorer work with non-linear function approximations? The code is available at Github https://github.com/arushijain94/ExplorationofGVFs.
@@
We benchmark GVFExplorer against several different baselines: (1) RoundRobin: uses a round-robin strategy to episodically sample from all target policies and uses off-policy Expected Sarsa for parallel estimation of multiple GVFs; (2) MixturePolicy: similar to RoundRobin, but uses a combined policy from all the given target policies; (3) SR: a Successor Representation (SR) based method which uses a intrinsic reward of total change in SR and reward weights to learn a behavior policy, as in(McLeod et al., 2021). (4) BPS: behavior policy search method originally designed for single target policy evaluation using a REINFORCE-style variance estimator (Hanna et al., 2017); we adapted it by averaging the variance across multiple GVFs (as in our objective). BPS results are limited to tabular settings due to scalability and performance issues. (5) UniformPolicy: a uniform sampling policy over the action space. We include all the relevant hyperparameters and implementation details in Section B.1.
@@
To answer the questions presented above, we consider different settings: (Setting 1): In a tabular setting, we examine two GVFs with distinct target policies but identical cumulants, $(\pi_{1},c),(\pi_{2},c)$. (Setting 2): In the same environment, we assess two GVFs with different target policy-cumulant pairs, $(\pi_{1},c_{1}),(\pi_{2},c_{2})$. (Setting 3): To verify the scalability of proposed method with high number of GVFs, we evaluate combinations of $4$ different target policies $\pi_{1}\dots\pi_{4}$ with $10$ different cumulants $c_{1}\dots c_{10}$, resulting in 40 GVFs. (Setting 4): In a continuous state environment with non-linear function approximator, we evaluate two distinct GVFs, $(\pi_{1},c_{1}),(\pi_{2},c_{2})$. Across these varied settings, we measure the MSE as a function of the the number of actions taken by the agent.
@@ 8.1 Tabular Experiments
We consider a $20\times 20$ grid with discrete states and four actions which moves the agent in four cardinal directions. The discount factor is $\gamma=0.99$. The environment is stochastic, where with $0.1$ probability, agent uniformly moves in any direction. In (Setting 1), the cumulant is placed at the top left corner with a reward of $\mathcal{N}(\mu=100,\sigma=5)$ and zero everywhere else. The agent starts from any non-goal state. The episode terminates either after $500$ steps or upon reaching the goal. Detailed description of the two target policies is provided in Section B.1. The true value function is calculated analytically as $V_{\pi}=(I-\gamma P_{\pi})^{-1}c_{\pi}$ for MSE computation. Figure 1 compares the averaged MSEs and the individual MSE for the estimated $\hat{v}_{\pi_{1}}^{c},\hat{v}_{\pi_{2}}^{c}$ for all algorithms. The results indicate that GVFExplorer reduces the MSE faster, indicating more precise value estimation given the same number of agent-environment interactions.
@@
(b) $\text{MSE}_{1},\text{MSE}_{2}$
Figure 1: Setting 1 in 20x20 Grid: MSE in value prediction for 2 target policies with the same cumulant $(\pi_{1},c),(\pi_{2},c)$ averaged over $25$ runs. We show results for baselines – RoundRobin, MixturePolicy, UniformPolicy, SR, BPS and GVFExplorer. (a) Averaged MSE as $\mathbb{E}_{i}[(V_{\pi_{i}}^{c}(s)-\hat{V}_{\pi_{i}}^{c}(s))^{2}]$. (b) $\text{MSE}_{1}:\mathbb{E}_{s}[(V_{\pi_{1}}^{c}(s)-\hat{V}_{\pi_{1}}^{c}(s))^{2}]$, and $\text{MSE}_{2}:\mathbb{E}_{s}[(V_{\pi_{2}}^{c}(s)-\hat{V}_{\pi_{2}}^{c}(s))^{2}]$. GVFExplorer shows lower MSE compared to other baselines.
Next, in (Setting 2), we evaluate two GVFs, each with a different target policy-cumulant pairs. For GVF1, the cumulant $c_{1}=\mathcal{N}(\mu=100,\sigma=5)$ is assigned to the top-left corner with zero elsewhere, while GVF2 receives $c_{2}=\mathcal{N}(\mu=50,\sigma=5)$ at the top-right corner only. Figure 2 presents the comparison of MSEs for different algorithms; GVFExplorer produces better value estimates compared to the baselines. We also qualitatively analyse the results, by comparing baseline RoundRobin and GVFExplorer. In Figure 2(b,e) the plots depict the averaged absolute difference between true and estimated GVF values across states, $\mathbb{E}_{i}[|V_{\pi_{i}}^{c_{i}}-\hat{V}_{\pi_{i}}^{c_{i}}|]$. GVFExplorer leads to smaller errors (duller colors), especially around the high variance goal regions. In Figure 2(c,f) we show GVFExplorer’s estimated variance for GVF1 and GVF2, demonstrating high variance regions. This plot highlights the need for strategic non-uniform sampling, by focusing on high variance regions to minimize environmental interactions. These variance plots show log-scale empirical values; most areas appear black, due to their relatively small magnitude compared to high variance regions.
(a) Average MSE
@@
(c) Var $\text{GVF}_{1}(\texttt{GVFExplorer})$
(d) $\text{MSE}_{1},\text{MSE}_{2}$
(e) Value Error(GVFExplorer)
(f) Var $\text{GVF}_{2}(\texttt{GVFExplorer})$
Figure 2: Setting 2 in 20x20 Grid: Evaluate two distinct GVF targets $(\pi_{1},c_{1})$ and $(\pi_{2},c_{2})$ averaged over 25 runs. We compare baselines – RoundRobin, MixturePolicy, UniformPolicy, SR, BPS with GVFExplorer. Green dots show GVF goals. (a) Averaged MSE, (d) $\text{MSE}_{1},\text{MSE}_{1}$. (b,e) Depicts averaged absolute error in GVFs value predictions for baseline RoundRobin and GVFExplorer. GVFExplorer demonstrate notably lower errors. (c,f) Shows estimated variance $\hat{M}^{c_{1}}_{\pi_{1}},\hat{M}^{c_{2}}_{\pi_{2}}$ in GVFExplorer, highlighting benefits of variance-based sampling. The color bar uses log scale, where vibrant colors indicate higher values.
Figure 3: Setting 3 in 20x20 Grid: 40 different GVFs comprising of all combinations of $4$ target policies and $10$ different cumulants. We show the averaged MSE over these GVFs for different baselines and GVFExplorer, where performance is averaged over $25$ runs. GVFExplorer method scales well to a large number of GVFs.
In the Setting 3, we evaluate our method’s ability to handle a large number of GVFs. We examine four target policies ($\pi_{n\in{1\dots 4}}$), each aligned with a cardinal direction, and ten cumulants ($c_{m\in{1\dots 10}}$), aiming to predict $40$ GVF combinations ($v_{\pi_{1\dots 4}}^{c_{1}}\dots v_{\pi_{1\dots 4}}^{c_{10}}$). Each GVF is associated with a state space region (“goal”), uniformly sampled and assigned cumulant value ranging from $50$ to $100$. The agent receives a zero cumulant signal on non-goal states. In Figure 3, we present a comparison of the average MSE across the various GVFs, illustrating the scalability of the proposed algorithm when increasing the number of GVFs. Notably, baseline SR shows limited scalability, because the varying cumulant scales affect the intrinsic behavior rewards (summation of SR weight change and reward weights). We omit BPS from this analysis due to its suboptimal performance in earlier experiments.

### v2

@@ 4.2 Theoretical Solution
The proof is presented in Sec. A.1. ∎
Theorem 4.1 describes how to iteratively update the behavior policy $\mu_{k}$ by using the return variance $M_{\pi_{i}}^{\mu_{k}}$. The policy $\mu_{k+1}$ selects actions proportional to their variance, meaning high-variance return $(s,a)$ pairs are explored frequently. By visiting high-variance return pairs, policy gains informative samples and reduce the overall uncertainty. Consequently, this process improves the GVF value predictions and decrease the number of interactions needed for effective learning.
Next Theorem 4.2 shows that each behavior policy update either decreases or maintains the overall variance across GVFs, ensuring consistent progress towards minimizing variance without oscillation. In simple terms, each policy update ensures either stability or improvement in data collection and value estimation efficiency.
@@ 5 Variance Function
The theorems provided in the previous section rely on the variance function $M_{\pi_{i}}^{\mu_{k}}$. Here, we study this variance function in detail.
@@
In an off-policy context [Sherstan et al. (2018), Jain et al. (2021)], introduced the variance function $M$, which estimates the variance in return under a target policy $\pi$ using data from a different behavior policy $\mu$. We will directly use this function $M$ as our variance estimator and present it here for completeness. The function $M$ for a state-action pair under $\pi$, with an importance sampling correction factor $\rho_{t}=\frac{\pi(a_{t}|s_{t})}{\mu(a_{t}|s_{t})}$, is defined as:
$M_{\pi}^{\mu}(s,a)$ $=\text{Var}_{a\sim\mu}\left(G_{t,\pi}|s_{t}=s,a_{t}=a\right)=\mathbb{E}_{a\sim\mu}\left[\delta_{t}^{2}+\gamma^{2}\rho_{t+1}^{2}M_{\pi}^{\mu}(s_{t+1},a_{t+1})|s_{t}=s,a_{t}=a\right]$ (4)
Here, $G_{t,\pi}$ is the return at time $t$, and $\delta_{t}=r_{t}+\gamma\mathbb{E}_{a^{\prime}\sim\pi}[Q_{\pi}(s_{t+1},a^{\prime})]-Q_{\pi}(s_{t},a_{t})$ is the TD error. We use Expected Sarsa (Sutton & Barto, 2018) to compute $\delta_{t}$, eliminating the need for IS by using the expected value of the next state-action pair under $\pi$ for bootstrapping, thus stabilizing the update and lowering the variance. $M_{\pi}^{\mu}(s,a)$ relates the variance under $\pi$ from the current state-action pair to the next, when sampled from $\mu$. This allows effective bootstrapping and iterative update using TD style method. The state variance function is defined as $M^{\mu}_{\pi}(s)=\sum_{a}\mu(a|s)\rho^{2}(s,a)M^{\mu}_{\pi}(s,a)$.
Note, true $Q_{\pi}$ is required to compute the TD error $\delta_{t}$ in Eq. 4. Following Sherstan et al. (2018), we substitute the value estimate $\hat{Q}$ for the true function $Q_{\pi}$ to compute $\delta_{t}$ in Eq. 4. Additionally, we use variance estimates $\hat{M}^{\mu_{k}}_{\pi}$ to update the next step policy $\mu_{k+1}$ instead of true variance in Eq. 3. This approach is similar to generalized policy iteration (Sutton & Barto, 2018), which simultaneously updates value estimator and improves the policy.
Next, we prove the existence of $M$ in Lemma 5.1, which was not covered in Jain et al. (2021). This proof establishes a loose upper bound on the IS ratio $\rho$, limiting the divergence of the behavior policy from the target policy for effective off-policy variance estimation. This aligns with methods like TRPO (Schulman et al., 2015) and Retrace (Munos et al., 2016), which stabilize policy updates by controlling divergence.
@@
(Variance Function $M$ Existence:) Given a discount factor $0<\gamma\leq 1$, the variance function $M$ exists, if the below condition satisfies, $\mathbb{E}_{a\sim\mu}\left[\rho^{2}(s,a)\right]<\frac{1}{\gamma^{2}}$ for all states.
@@
Proof in Sec. A.2. ∎
Note, the objective in Eq. 2, might violate the above constraint on $\rho$; we empirically clip $\rho$ to mitigate this problem. Additionally, IS requires $\mu(a|s)=0$ when $\pi(a|s)=0$. We empirically ensure $\mu(a|s)>\varepsilon<<1$ for all actions. The same constraint is added for all the baselines for fair comparison.
@@ 6 Algorithm
We present GVFExplorer algorithm, detailed in Algorithm 1. Our approach uses two networks: $Q_{\theta}$ for value function and $M_{w}$ for variance, each with $\mathcal{N}$ heads (one head for each GVF). Starting with a randomly initialized behavior policy, the agent observes cumulants for $\mathcal{N}$ GVFs at each step and updates $Q_{\theta}$ using off-policy TD. We use Expected Sarsa (Sutton & Barto, 2018) for both $Q$ and $M$, eliminating off-policy corrections. The target $Q$ updates follow:
$Q_{tar}(s_{t},a_{t},s_{t+1})=c_{t}+\gamma\sum_{a^{\prime}}\pi(a^{\prime}|s_{t+1})Q_{\theta}(s_{t+1},a^{\prime}).$ (5)
We use the TD error from $Q$ learning, $\delta_{Q}=Q_{tar}-Q_{\theta}$, to update target $M$,
$M_{tar}(s_{t},a_{t},s_{t+1})=\delta_{Q}^{2}+\gamma^{2}\sum_{a^{\prime}}\pi(a^{\prime}|s_{t+1})M_{w}(s_{t+1},a^{\prime}).$ (6)
Both networks are updated via MSE loss. The behavior policy is iteratively updated using the new variance estimates for $K$ steps, with learned $Q$ values used for MSE metrics in Eq. 1.
To ensure reliable estimates, we initialize $M$ values to small non-zero constants and apply epsilon exploration, which decays over time, ensuring coverage of the state-action space. This guarantees that agents visit a broad range of state-action pairs early on, preventing issues of zero variance for unvisited pairs. We applied epsilon-exploration to both GVFExplorer and the baselines for fair comparison.
We also use techniques like Experience Replay Buffer for data reuse and target networks for both $Q$ and $M$ to improve learning stability. Expected Sarsa is used consistently across all baselines for fair comparison. Refer Algorithm 1 for further details.
Input: Target policies $\pi_{i\in\{1,\dots n\}}$, initialized behavior policy $\mu_{1}$, replay buffer $\mathcal{D}$, primary networks $Q_{\theta},M_{w}$ (small non-zero $M$), target networks $Q_{\bar{\theta}},M_{\bar{w}}$, learning rates $\alpha_{Q}$, $\alpha_{M}$, mini-batch size $b$, trajectory length $T$, target update frequency $l=100$, value/variance update frequencies $p=4$, $m=8$, training steps $K$, exploration rates $\varepsilon_{0}$, $\varepsilon_{\text{decay}}$, $\varepsilon_{\min}$
@@
2for environment step $k=1,\dots K$ do
3       Set exploration rate: $\varepsilon_{k}=\max(\varepsilon_{\min},\varepsilon_{0}\cdot\varepsilon_{\text{decay}}^{k})$
4       Select action $a_{t}\sim\varepsilon$-greedy policy $\mu_{k}(\cdot|s_{t})$
5       Observe next state $s_{t+1}$ and cumulants $c_{t}=\textit{Vector(size}(n))$
6       Store transition $(s_{t},a_{t},s_{t+1},c_{t})$ in $\mathcal{D}$
7
@@ 6       Store transition $(s_{t},a_{t},s_{t+1},c_{t})$ in $\mathcal{D}$
9             //Update the Value $Q_{\theta}$ network
@@
11             Update $Q_{\theta}$ using MSE loss $(Q_{tar}(s_{t},a_{t})-Q_{\theta}(s_{t},a_{t}))^{2}$, where $Q_{tar}$ is Eq. 5.
@@ 17             Sample mini-batch of size $b$ of transition $(s_{t},a_{t},s_{t+1},c_{t})\sim\mathcal{D}$
18             Update $M_{w}$ using MSE loss $(M_{tar}(s_{t},a_{t})-M_{w}(s_{t},a_{t}))^{2}$, where $M_{tar}$ is Eq. 6.
19
20       end if
21
@@
23             $\bar{w}=w$ and $\bar{\theta}=\theta$  //Update both target networks weights
@@
30Returns GVFs Values $V_{i}(s)=\sum_{a}\pi_{i}(a|s)Q_{\theta}^{i}(s,a)$ for $i=\{1,\dots,n\}$
@@ 7 Experiments
We investigate the empirical utility of our proposed algorithm in both discrete and continuous state environments. Our experiments are designed to answer the following questions: (a) How does GVFExplorer compare with the different baselines (explained below) in terms of convergence speed and estimation quality? (b) Can GVFExplorer handle a large number of GVFs evaluations? (c) Can GVFExplorer work with non-stationary GVFs which change with time? (d) Can GVFExplorer work with non-linear function approximations and complex Mujoco environments?  [^1] The code is available at Github.
@@
We use Off-policy Expected Sarsa updates for parallel GVF estimations for all the experiments (including baselines) for fair comparison. We benchmark against several different baselines: (1) RoundRobin: uses a round-robin strategy sampling episodically from all target policies (2) MixturePolicy: Aggregated policy sampling from all target policies; (3) SR: a Successor Representation (SR) method using intrinsic reward of total change in SR and reward weights to learn behavior policy (McLeod et al., 2021). (4) BPS: behavior policy search method originally designed for single policy evaluation using a REINFORCE variance estimator (Hanna et al., 2017); we adapted it by averaging variance across multiple GVFs (similar to our objective). BPS results are limited to tabular settings due to scalability issues with it. (5) UniformPolicy: a uniform sampling policy over the action space. Implementation details and hyperparameters are in App. B.
Type of Cumulants.
We experiment with three different types of cumulants, similar to McLeod et al. (2021)constant with a fixed value; distractor, a stationary signal with fixed mean and constant variance (normal distribution); drifter, a non-stationary cumulant with zero-mean random walk with low variance (vary with time). Further description of cumulants is in Sec. B.2.
@@
To answer the questions presented above, we consider different settings: (Two Distinct Policies & Identical Cumulants): In a tabular setting, we examine two GVFs with distinct target policies but identical distractor cumulant, $(\pi_{1},c),(\pi_{2},c)$. (Two Distinct Policies & Distinct Cumulants): In the same environment, we assess two GVFs with distinct target policy and distinct distractor cumulant with different fixed means, $(\pi_{1},c_{1}),(\pi_{2},c_{2})$. (Large Scale Evaluation with $40$ distinct GVFs): To verify the scalability of proposed method with high number of GVFs, we evaluate combinations of $4$ different target policies $\pi_{1}\dots\pi_{4}$ with $10$ different constant cumulants $c_{1}\dots c_{10}$, resulting in 40 GVFs. (Non-Stationary Cumulants in FourRooms): In FourRooms environment, we assess with two distinct GVFs - stationary distractor and non-stationary drifter cumulant$(\pi_{1},c_{1}),(\pi_{2},c_{2})$. (Non-Linear Function Approximation): In a continuous state environment with non-linear function approximator, we evaluate two distinct distractor GVFs, $(\pi_{1},c_{1}),(\pi_{2},c_{2})$. (Mujoco environments): In Mujoco environments – walker and cheetah – evaluate different GVF tasks like walk, run and flip. Across these varied settings, we measure the averaged MSE across multiple GVFs.
@@ 7.1 Tabular Experiments
We conducted experiments in $20\times 20$ gridworld with four cardinal actions and a tabular $20\times 20$ FourRooms environment for added complexity. The discount factor is $\gamma=0.99$, and the environment is stochastic with a $0.1$ probability of random movement. The cumulants are zero everywhere except for at the goals. Episode terminates after $500$ steps or upon reaching the goal. True value function for MSE computation is calculated analytically $V_{\pi}=(I-\gamma P_{\pi})^{-1}c_{\pi}$. Detailed description of target policies and cumulants is provided in Sec. B.3. Table 1 summarizes the below results for tabular experiments.
In Two Distinct Policies & Identical Cumulants, we consider gridworld environment with distractor cumulant at top left corner with a reward drawn from normal distribution. Fig. 1(a) shows the averaged MSE across the two GVFs, with GVFExplorer showing much lower MSE compared to baselines.
Next, in Two Distinct Policies & Distinct Cumulants, we consider two distinct distractor cumulant (with different mean) GVFs placed at top-left and top-right corner respectively. Fig. 2(a) shows GVFExplorer with reduced MSEs compared to baselines. Figs. 2(b) and 2(c) qualitatively analyze the average absolute difference between true and estimated GVF values across states, $\mathbb{E}_{i}[|V_{\pi_{i}}^{c_{i}}-\hat{V}_{\pi_{i}}^{c_{i}}|]$, showing smaller errors (duller colors) for GVFExplorer. Fig. 8 (in Sec. B.3.2) presents the individual variance and MSE for both GVFs in GVFExplorer. Further, we conduct an ablation study to experiment with how GVFExplorer performance changes with poorer feature approximations. Fig. 10 (in Sec. B.3.3) shows that MSE increases as the feature quality deteriorates, but GVFExplorer remains robust with moderately coarse approximations.
For Non-Stationary Cumulant in FourRooms, we evaluate the performance in FourRooms (FR) environment (Sutton et al., 1999) with two distinct GVFs: stationary distractor cumulant and a non-stationary drifter cumulant which changes value over time. As shown in Fig. 1(b), GVFExplorer reduces MSE faster than other baselines, even with the non-stationary cumulant. Fig. 11 (in Sec. B.3.4) demonstrates the effectiveness of GVFExplorer in tracking the non-stationary cumulant signal in the later stages of learning.
In Large Scale Evaluation with 40 Distinct GVFs, we evaluate our method’s scalability to large number of GVFs (refer Sec. B.3.5). We use constant cumulants with values ranging in $[50,100]$. Fig. 1(c) compares the average MSE across the GVFs, showing that GVFExplorer scales well with an increasing number of GVFs. In contrast, the SR baseline struggles with scalability due to the varying cumulant scales affecting the intrinsic reward (the summation of all SRs and reward weights) of behavior policy.
(a) Two Distinct Policies & Identical Cumulants
(b) Non-Stationary Cumulant in FourRooms
(c) Large Scale Evaluation with 40 Distinct GVFs
Figure 1: MSE Performance: Averaged MSE over $25$ runs with standard error in different experimental settings. GVFExplorer demonstrate notably lower MSE compared to the baselines.
@@
(c) Value Error (GVFExplorer)
Figure 2: Two Distinct Policies & Distinct Cumulants: Evaluate averaged MSE over 25 runs with two distinct distractor GVFs $(\pi_{1},c_{1}),(\pi_{2},c_{2})$ in gridworld . Green dots at top show two GVF goals. (a) Averaged MSE, (b) averaged absolute error in GVFs value predictions for baseline RoundRobin and (c) GVFExplorer. The color bar uses log scale & vibrant colors indicate higher values.